OpenStudy (laronjaxon):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3? y = –x – 1 y = x + 1 y = x + 5 y = –x + 5

OpenStudy (anonymous):

I will guide you will be able to get the solution. For perpendicular lines, the product of their slopes = -1. So first get the slope of the line you looking for, how much do you get?

OpenStudy (laronjaxon):

where do i add the -1?

OpenStudy (laronjaxon):

i solved it nvm thank u tho

OpenStudy (anonymous):

Say u have 2 lines perpendicular to each other. y = ax + c1 and y = bx + c2. We say a*b = -1. In our case you given 1 line with slope -1 therefore say b*-1 = -1 where b is the slope of the line you looking for, therefore our slope is 1 (positive) u see that?

OpenStudy (anonymous):

you welcome